Answer
$0.19.$
Work Step by Step
In a probability model the sum of all probabilities equals $1$.
Hence $0.03+P(\text{18-20 years old})+0.23+0.29+0.25+0.01=1\\P(\text{18-20 years old})=1-(0.03+0.23+0.29+0.25+0.01)=0.19.$
You need to log in to continue
You can help us out by revising, improving and updating this answer.
Update this answerAfter you claim an answer you’ll have 24 hours to send in a draft. An editor will review the submission and either publish your submission or provide feedback.